5 mol of an ideal gas at a constant temperature if 0°c. If 2kj of energy is inputted into the system and the net change of internal energy ∆u=0, by what factor does the pressure of the gas change.

1
Expert's answer
2022-02-24T09:05:48-0500

"A=\nu RT\ln\frac{p_2}{p_1},"

"\frac{p_2}{p_1}=e^{\frac{A}{\nu RT}}=1.2."